OpenStudy (anonymous):

Here's one more that I need the steps to not the answer, If I can get the steps I can get the answer, thanks. Given the quadratic function f(x) = x2 – 12x + 36, find a value of x such that f(x) = 9.

OpenStudy (lgbasallote):

f(x) = x^2 - 12x + 36 AND f(x) = 9 since f(x) = f(x) we can therefore say that x^2 - 12x + 36 = 9

OpenStudy (anonymous):

Thank you. I just factor from there to solve correct?

OpenStudy (lgbasallote):

subtract 9 from both sides then yes you can factor it out afterwards

OpenStudy (anonymous):

K, thank you so much :)

OpenStudy (anonymous):

Is this correct, it doesn't look like it is to me? X^2 – 12x + 36 = 9 X^2 – 12x + 36 – 9 = 9 – 9 X^2 – 12x + 27 X^2 – 6 * x * 2 + 27 (x – 2)^2= 2

OpenStudy (lgbasallote):

X^2 – 6 * x * 2 + 27 (x – 2)^2= 2 what did you do there?

OpenStudy (anonymous):

Not sure what I was suppose to do, the instuctions from the instrutor said to find the square root of both sides. But I don't know how to find the square root of x

OpenStudy (lgbasallote):

completing the square?? <---that is another method in finding x..what im doing is factoring

OpenStudy (anonymous):

This is the example she gave us to go off of, she wants the problem done like this. Given the quadratic function f(x) = x2 – 8x + 16, find a value of x such that f(x) = 4 Hint: First, the given function can be factored using perfect square trinomials, and then set the function equal to the value given for f(x). Look at example 5 on p. 706 for a similar example. Answer: x = 6 and x = 2 Explanation: Set x2 – 8x + 16 = 4 and factor x2 – 8x + 16 by using the rule for factoring perfect square trinomials to get: (x – 4)2 = 4 Find the square root of both sides to get: x – 4 = ± 2 Solve for x to get: x = 4 + 2 = 6 and 4 – 2 = 2

OpenStudy (lgbasallote):

lol why didnt you say so....it would be done IN THE BEGINNING...okay let's start again... x^2 - 12x + 36 = 9 factor out x^2 - 12x + 36

OpenStudy (anonymous):

Um, ok

OpenStudy (anonymous):

That would be using the perfect square rule correct? That's what I followed when I facotered it out earlier

OpenStudy (lgbasallote):

yup it is...and btw...x^2 - 12x + 27 isnt a perfect square trinomial that's why we couldnt do it there

OpenStudy (anonymous):

K, so I would follow the same rules that I did earlier to get x^2 - 12x - 36 (x - 2)^2 still following what the book says, which confuses me even more

OpenStudy (lgbasallote):

(x-2)^2 = x^2 - 4x + 4! (x-2)^2 is not a fixed for every thing that looks like this :p

OpenStudy (anonymous):

I was just following the book and it skips around a lot :P So after that step is when the square root comes in?

OpenStudy (lgbasallote):

(x-2)^2 is not even right yet...let's factor out x^2 - 12x + 36 first

OpenStudy (lgbasallote):

i assume you have an idea on factoring right?

OpenStudy (anonymous):

Yeah, I just hate doing factoring so I always forget to do that part :)

OpenStudy (lgbasallote):

rewrite x^2 - 12x + 36 as \[x^2 - 6x - 6x + 36\] now facot that

OpenStudy (anonymous):

x^2 - 12x + 36 x^2=(x*x) 12x= (4*x*4) 36= (6*6)

OpenStudy (anonymous):

-6x(x^2 +36)

OpenStudy (lgbasallote):

what are you doing? hahaha :))) you factor by the one i stated \[x^2 - 6x - 6x + 36\]

OpenStudy (anonymous):

That's how my brain works now, very frustrating :)

OpenStudy (lgbasallote):

group it as \[(x^2 - 6x) + (-6x + 36)\] then factor out common factors

OpenStudy (lgbasallote):

btw..6x(x^2 + 36) is equal to 6x^3 + 216x haha

OpenStudy (anonymous):

That's what I hate about my brain I know that step but it passes it completely :) (x^2 - 6x)+(-6x + 36) -6x(x^2 + 36) Then I believe you would factor out the x^2, at least that's what she had us do last class.

OpenStudy (lgbasallote):

lol =)))) i now see what you're doing wrong! let me first ask you...do you know what "factor" means?

OpenStudy (anonymous):

Used to mean to break the term down to prime numbers but my instructor doesn't like that. She showed us one way but for the tests we're expested to know exactly what she wants without telling us.

OpenStudy (lgbasallote):

"factors" means terms that are multiplied...im talking about the basic-est math hehe

OpenStudy (anonymous):

Ah ok, why couldn't she say that, it would make things so much easier :) So after factoring the 36 to 6*6 I would take the steps I was jumping?

OpenStudy (lgbasallote):

hmmm....let me do one step and maybe you can do the other....ill factor out x^2 - 6x watch what i do (x)(x) <===that's x^2 (6)(x) <=== 6x x^2 - 6x = (x)(x) - (6)(x) now see how both of them have an x? that means i can "factor" out x x(x - 6) <---factored form of x^2 - 6x

OpenStudy (lgbasallote):

now do -6x + 36

OpenStudy (anonymous):

-6x= 3x*2 36= 3*12 3x(2+12)

OpenStudy (lgbasallote):

btw what i did was continuation of \((x^2 - 6x) + (-6x + 36)\) i simply factor out each "group"

OpenStudy (anonymous):

Or would the x be with the 2, that part always throws me off

OpenStudy (lgbasallote):

-6x = (-3)(x)(2) 36 = (3)(3)(2)(2) ^see how 36 doesnt have an x? therefore you cant factor out ..try again

OpenStudy (lgbasallote):

therefore you cant factor out x is what i meant*

OpenStudy (anonymous):

So it would be -3(2x + (3)(2)(2) or would the outside nuber be 3. The teacher never covered this in class at all and the example in the book don't show it

OpenStudy (lgbasallote):

then ill be teaching you something new :P 2 is still common between the two...factor it out

OpenStudy (anonymous):

So it's actually 2(3x + 3*3*2)

OpenStudy (lgbasallote):

noooo!! 3 AND 2 are BOTH common

OpenStudy (anonymous):

so I factor out both at the same time?

OpenStudy (lgbasallote):

yes!

OpenStudy (anonymous):

3*2(x+3*2), correct?

OpenStudy (lgbasallote):

almost.... remember how -6x was NEGATIVE?

OpenStudy (anonymous):

Crap I forgot that LOL. -3*2(x + 3*2)

OpenStudy (lgbasallote):

hmmm...seems i need to teach you what happens in factoring....i'll use the first factor...x^2 - 6x when i factored out x, this is what happened \[\large x^2 - 6x = x(\frac{x^2}{x} - \frac{6x}{x}) =x(\frac{x^{\cancel{2}}}{\cancel{x}} - \frac{6\cancel{x}}{\cancel{x}}) = x(x - 6)\]

OpenStudy (lgbasallote):

do you get my point?

OpenStudy (anonymous):

yeah, that part I get. The part I don't get is factoring with only one negative number. I don't think you can make x a negative.

OpenStudy (lgbasallote):

now what we did in -6x + 36...what did you factor out again?

OpenStudy (anonymous):

I actored out the 3 and the 2

OpenStudy (lgbasallote):

NEGATIVE 3 and 2 :P wwhich is also equal to -6 agree?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

so \[-6x + 36 = -6(\frac{-6x}{-6} + \frac{36}{-6})\] what's the answer there?

OpenStudy (anonymous):

-6(1x - 6)

OpenStudy (anonymous):

Duh I see now :)

OpenStudy (lgbasallote):

YAYY!!!

OpenStudy (lgbasallote):

so the equation becomes \[x(x - 6) - 6(x - 6) = 9\] factor again

OpenStudy (anonymous):

Would that be 3x(x+3) or is that completely screwed up, sorry I have a hard time going over thing that are new. Even if i just went over it :)

OpenStudy (lgbasallote):

hmm think of (x-6) as a ax - 6a..how would you factor that?

OpenStudy (anonymous):

a(x-3*3) I think

OpenStudy (lgbasallote):

a(x-6) is the right one :P now substitute a back

OpenStudy (anonymous):

ax-6a if I multiply it

OpenStudy (lgbasallote):

then you just went back to the original equation we had! i meant change a to (x-6) again...wasnt that what a meant?

OpenStudy (anonymous):

OK, (x-6)(x-6)

OpenStudy (lgbasallote):

GOOD! and that is...?

OpenStudy (anonymous):

x^2 - 6x - 6x +36 x^2 +12x +36

OpenStudy (lgbasallote):

do you notice that you just went back to what you had a while ago :P

OpenStudy (anonymous):

I feel like such an idiot, I knew that :)

OpenStudy (lgbasallote):

Hint: \(a \times a = a^2\)

OpenStudy (anonymous):

So did that solve the problem or is there more to it?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

I thought there was a part that required square root or was it done already

OpenStudy (lgbasallote):

i dont know what you're talking about :P haha you have \((x-6)(x-6)\) i am telling you that \(a \times a = a^2\)

OpenStudy (anonymous):

Ok, I'm confused what does a*a=a^2 mean. I knew that much :)

OpenStudy (lgbasallote):

it means that when you multiply two terms that are the same say \((x-6) \times (x-6)\) that is equal to \((x-6)^2\)

OpenStudy (anonymous):

Haha, I knew that :D

OpenStudy (lgbasallote):

so now we have \[(x-6)^2 = 9\] time to do your square root thing :PPP

OpenStudy (anonymous):

Oh lord, I have never fully understood that part :)

OpenStudy (lgbasallote):

you take the square root of both sides

OpenStudy (anonymous):

Stupid question, how do you get the square root of a variable? I know the square root of 9 is 3

OpenStudy (lgbasallote):

hmm let me ask this to you...what do you think is square root? what does it mean?

OpenStudy (anonymous):

What times itself equals a number

OpenStudy (anonymous):

Like 3*3=9

OpenStudy (lgbasallote):

right! in other words what number when squared will result to that number correct?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

square root of 9 means what number do you need to square to get 9...so what do you square to get (x-6)^2?

OpenStudy (lgbasallote):

Hint: \(\sqrt{a^2} = a\)

OpenStudy (anonymous):

would I make it x^2 - 36 before I find the square root? Sorry, my kids are trying to figure it out the wrong ways and confusing me in the process :)

OpenStudy (lgbasallote):

noooo look at the hint i wrote...imagine x - 6 is a

OpenStudy (anonymous):

\[\sqrt{a-6}^{2}=a + 36\] correct? or would it be different

OpenStudy (lgbasallote):

nooo where did you get that 36?

OpenStudy (lgbasallote):

and a....what?? that's x not a

OpenStudy (anonymous):

LOL -6 * -6

OpenStudy (lgbasallote):

why....why did you square it -_-

OpenStudy (anonymous):

Told you I suck at this part :)

OpenStudy (lgbasallote):

just look at the one i wrote square root of a^2 is a...now..what's square root of (x-6)^2??